# Quantum control of continuous systems via nonharmonic potential modulation

### v3

@@ 3.2 Unitaries
$\hat{P}$ is a projector onto a subspace, $\hat{U}_{\text{T}}$ is a target unitary, and $\hat{U}$ is a unitary generated through Eq. (9). Taking advantage of the optimized position of the potential $u(\tau)$ and additional slight optimized modulation of the depth $a(\tau)$, we present examples of unitaries, e.g., $\sigma_{x}$ or Hadamard operations, for above-mentioned subspaces. In Fig. 3(a,b) we show an exemplary case of four-legged-cat basis and $\sigma_{x}$ operation performed with a Gaussian potential, where the basis states are given by
@@
$\mathcal{N}_{\pm}$ are the normalization constants, and we use $\beta=2$. Again, the protocol is performed with high fidelity, $\mathcal{F}_{\hat{U}}\approx 99\%$. In Fig. 4 we show additional examples for Gaussian potential, including, among others, GKP basis, where the basis is spanned by two mutually displaced GKP states,
@@
where $d_{3}$ is defined as in (22) and the squeezing and displacement need to be enough for the states not to overlap. Further unitary implementations, but for the cosine potential with the constraint (2.1) are presented in the Appendix A.
Figure 3: (a,b) Two orthogonal four-legged-cat states spanning a two-level subspace evolved in a position- and depth-controlled single Gaussian well, realizing a $\sigma_{x}$ operation. (c) Cat state preparation using a double-well optimal control. The superscript L(R) signifies the state centered in the left(right) potential well. (d,e) State transfer between the wells—each of the vibrational states, $\psi_{0}(x)$ and $\psi_{1}(x)$, is transferred to the other well, without altering the relative phase.
@@ 3.3 Double-well potential landscape
$\hat{M}^{\prime}=\begin{pmatrix}\innerproduct{\psi^{\text{L}}_{+}(x,\tau_{\text{max}})}{\psi^{\text{R}}_{+}(x)}&\innerproduct{\psi^{\text{L}}_{+}(x,\tau_{\text{max}})}{\psi^{\text{R}}_{-}(x)}\\ \innerproduct{\psi^{\text{L}}_{-}(x,\tau_{\text{max}})}{\psi^{\text{R}}_{+}(x)}&\innerproduct{\psi^{\text{L}}_{-}(x,\tau_{\text{max}})}{\psi^{\text{R}}_{-}(x)}\end{pmatrix},$ (36)
@@
as the reward function. In Fig. 3(d,e) we show the example of a state of two lowest vibrational levels of the left well transferred to the right one with $\mathcal{F}_{\text{st}}\approx 99\%$.
Figure 4: Evolution of selected orthogonal states in a Gaussian potential characterized by $\eta=0.25$ and with optimally controlled displacement $u(\tau)$ and intensity $a(\tau)$. The solid line shows $u(\tau)$ and the protocol lasts $\tau_{\text{max}}/2\pi=9$. (a-b) $\sigma_{x}$ unitary within a subspace spanned by $\psi_{0}$ and $\psi_{1}$ Fock states. The fidelity reads $\mathcal{F}_{\hat{U}}\approx 99.8\%$. (c-d) Hadamard unitary within a subspace spanned by GKP states with $r=0.7$, $(d_{1},d_{2},d_{3})=(-\sqrt{6\pi},0,\sqrt{6\pi})$. The fidelity reads $\mathcal{F}_{\hat{U}}\approx 96.4\%$. (e-f) $\sigma_{y}$ unitary within a subspace spanned by four-legged-cat states states with $\beta=2$. The fidelity reads $\mathcal{F}_{\hat{U}}\approx 96.2\%$.
@@ 3.4 Single-shot discrimination
where $k$ is chosen such that two phase-imprinted states are nonoverlapping in phase space. After the phase imprinting, a selective measurement takes place, which can be realized in, e.g., optical tweezer setup through the subsequent release of the trap in which time-of-flight evolution reveals spatially separated detection clicks for each of the states [112, 44]. The reward function for such a momentum-kick protocol involves a sum of an equally weighted fidelities,
$\mathcal{F}_{\text{mk}}=\frac{1}{2}\sum_{j=\pm}\left|\innerproduct{\psi^{j}(x,\tau_{\text{max}})}{\psi^{j}_{k}(x)}\right|^{2}$ (39)
Note that in contrast to the unitary implementation, here the final relative phase between the states does not matter, as we aim only to distinguish the initial states. As an example, in Fig. 5 $\psi^{\pm}(x)$ are taken to be the two lowest eigenstates of the well and the protocol is shown in the schematic way. In Fig. 6 we show in detail specific realizations of this and other momentum-kick protocols, utilizing a double-well Gaussian potential. They involve also other choices of orthogonal states for the discrimination, namely eigenstates of both $\sigma_{x}$ and $\sigma_{y}$ Pauli matrices. Such discrimination procedures then correspond to $\sigma_{x}$ and $\sigma_{y}$ measurements within a given two-level subspace.
The second discrimination protocol involves a double-well potential landscape and utilizes spatial separation instead of the momentum one (see Fig. 5). Initially, the states are fully confined to the left well and their discrimination is realized through a selective stealing protocol—if the state is $\psi^{+}(x)$, then it ends up in the second well after the optimal shaking, while $\psi^{-}(x)$ stays in the initial one. Then, large spatial separation between the wells enables selective measurement. The reward function for this protocol, the selective-stealing fidelity, reads
$\mathcal{F}_{\text{ss}}=$ $\frac{1}{2}\Big{[}\left|\innerproduct{\psi_{-}^{\text{L}}(x,\tau_{\text{max}})}{\psi_{-}^{\text{L}}(x)}\right|^{2}$
$\quad\quad+\left|\innerproduct{\psi_{+}^{\text{L}}(x,\tau_{\text{max}})}{\psi_{+}^{\text{R}}(x)}\right|^{2}\Big{]},$ (40)
where the definitions (34) and (35) are used. We show the exemplary case of distinguishing two lowest vibrational states in two top panels of Fig. 7, while further examples, including discrimination of $\sigma_{x}$ and $\sigma_{y}$ eigenstates, are presented in Appendix C.
@@ 3.5 Cooling
Figure 6: Examples of the implementation of momentum kick protocols via an optimally controlled single-well Gaussian potential with $\eta=0.25$, $\tau_{\text{max}}/2\pi=6$, and $k=2$. (a-b) $\psi^{+}(x)=\psi_{0}(x)$ and $\psi^{-}(x)=\psi_{1}(x)$, (c-d) $\psi^{\pm}(x)=[\psi_{0}(x)\pm\psi_{1}(x)]/\sqrt{2}$, and (e-f) $\psi^{\pm}(x)=[\psi_{0}(x)\pm i\psi_{1}(x)]/\sqrt{2}$. Momentum-kick fidelities read (a-b) $\mathcal{F}_{\text{mk}}\approx 98.5\%$, (c-d) $\mathcal{F}_{\text{mk}}\approx 99.4\%$, and (e-f) $\mathcal{F}_{\text{mk}}\approx 99.5\%$.
$\mathcal{F}_{\text{ss}}^{\text{cool}}=$ $\frac{1}{N+1}\Big{[}\left|\innerproduct{\psi_{0}^{\text{L}}(x,\tau_{\text{max}})}{\psi_{0}^{\text{L}}(x)}\right|^{2}$
$\quad\quad+\sum_{n=1}^{N}\left|\innerproduct{\psi_{n}^{\text{L}}(x,\tau_{\text{max}})}{\psi_{n}^{\text{R}}(x)}\right|^{2}\Big{]},$ (41)
@@
Here, $N$ characterizes the amount of excited states that are brought to the other well. The higher $N$, the more thermal state can be brought to the ground state. However, higher temperature also means a lower chance of success, as the presence of the ground state in the left well is conditioned on the subsequent destructive measurement in the right well. Hence, on average, the success rate of the protocol is proportional to the initial occupation of the ground state. Fig. 7 demonstrates this process for the four lowest states of the well, with the higher three states becoming spatially separated from the ground state. The method involving momentum-kick protocol is analogous, but with the excited fraction being kicked out of the well. The reward function is thus
$\mathcal{F}_{\text{mk}}^{\text{cool}}=$ $\frac{1}{N+1}\Big{[}\left|\innerproduct{\psi_{0}(x,\tau_{\text{max}})}{\psi_{0}(x)}\right|^{2}$
$\quad\quad+\sum_{n=1}^{N}\left|\innerproduct{\psi_{n}(x,\tau_{\text{max}})}{\psi_{n}^{k}(x)}\right|^{2}\Big{]},$ (43)
@@
3.6 State stabilization
Note that undriven evolution is nonlinear and hence, the prepared state gets distorted during the evolution. In contrast, when evolved in the perfectly harmonic potential, the state undergoes rotation in the phase space with the trap period. In a realistic setup, such a state preservation in the rotating frame is destroyed by the nonharmonicity of the potential. The remedy lies in either decreasing the nonharmonicity or, alternatively, mimicking perfect phase space rotation through optimal control [60]. We present the latter method, which consists of two optimization steps. The first one involves using the average fidelity with respect to the perfectly rotating state during a multiple of the trap period $\tau_{\text{max}}=n2\pi$ as a reward function,
Figure 7: The accelerated algorithmic cooling method by spatial separation of the ground and excited fractions realized through optimal control of two wells. The selective-stealing cooling fidelity reads $\mathcal{F}_{\text{ss}}^{\text{cool}}\approx 97\%.$
Figure 8: (a) Temporal evolution of a cat state in a perfectly harmonic potential. It rotates in the phase space and fully revives every trap period. (b) Evolution of a cat state in a tweezer characterized by $\eta=0.25$. (c) Evolution of a cat state in a tweezer characterized by $\eta=0.25$, however with optimally controlled potential position and depth. (d) Fidelities between a cat state rotating in a perfectly harmonic potential and in Gaussian potentials with different nonharmonicities without (shades of green) and with (red) optimal control.
[…]
where
|$\ket{\psi_{\text{rot}}(\tau)}=e^{-i\hat{a}^{\dagger}\hat{a}\tau}\ket{\psi(0)}$ (46)
$\mathcal{F}=\left|\innerproduct{\psi(x,\tau_{\text{max}})}{\psi(x,0)}\right|^{2}.$ (47)
The second step guarantees high fidelity, should the protocol be applied consecutively many times, while the first one ensures that state is not distorted at any time. In Fig. 8 we present an example of an optimally controlled cat state stabilization. The optimization is conducted for a single trap period, $\tau_{\text{max}}=2\pi$ with both position and depth control of a Gaussian potential. Subsequently, the optimized control is applied six times consecutively over six trap periods and compared with the bare evolution in a Gaussian potential. The results showcase that the presented optimal stabilization of the state surpasses that of a Gaussian potential with a relatively low nonharmonicity $\eta=0.05$, within this particular time scale.
@@ 4 Feasibility
After presenting several protocols, let us now address aspects relevant to their experimental implementation. From now on, we will focus on the Fock state preparation protocol with a single-well Gaussian potential as an exemplary case. In subsection 4.1 we analyze how fast the optimized protocol can be and how excited the system is during its evolution. Then, in subsection 4.2 we check the robustness of the state preparation in the presence of noise and decoherence modeled by the stochastic fluctuations of the potential. Finally, subsection 4.3 addresses the issue of satisfying one-dimensional approximation in the specific case of an atom in an optical tweezer.
@@ 4.1 Quantum speed limit
Concerning the protocol’s speed, a fundamental restriction exists, the quantum speed limit [113], which gives the minimal time that is needed to perform a specific state transfer with a given fidelity threshold. Well-understood for two-level systems driven by static Hamiltonians [113, 114, 115], quantum speed limits have also been analyzed for time-dependent multilevel systems [116, 117], and, specifically, for the ground to the first excited state transfer in a Bose-Einstein condensate [51]. We analyze a Gaussian potential case to find a minimum time $\tau_{\text{QSL}}$ needed to perform the excitation from the ground to the first excited state, $\psi_{0}(x)\rightarrow\psi_{1}(x)$, with some infidelity threshold for different values of nonharmonicity $\eta$ In the regime of interest, $\eta\sim 0.1-0.5$, there is no strong scaling of $\tau_{\text{QSL}}$ with the nonharmonicity and achievable fidelity scales exponentially with the protocol time. For the example of an optical tweezer, it implies that, without noise, the deeper the tweezer, the more bound states accessed and the faster the protocol (and hence requiring more control bandwidth), as $\omega\sim V_{0}^{1/2}$. We then analyze the number of excitations averaged throughout the protocol,
Figure 9: (a) Infidelity of the ground to the first excited state transfer, $\psi_{0}(x)\rightarrow\psi_{1}(x)$, for a Gaussian potential as a function of the length of protocol $\tau_{\text{max}}$ and nonharmonicity $\eta$. The scaling of the curve is as expected from the previous studies [51] and reveals that a quantum speed limit $\tau_{\text{QSL}}$ does not scale strongly with $\eta$. (b) Average excitation number as a function of $\tau_{\text{max}}$ and $\eta$. (c,d) The minimal achievable infidelity for given values of position ($\gamma_{1}$) and depth ($\gamma_{2}$) noises for a Gaussian potential. (e,f) The minimal achievable infidelity for given values of displacement $U_{\text{rms}}$ and intensity $A_{\text{rms}}$ noises for a rubidium-87 atom in an optical tweezer with the waist $w_{0}=710$nm. The value of nonharmonicity $\eta$ is tuned via the tweezer’s depth $V_{0}$.
[…]
as it is relevant for analytical quantum speed limits [115, 118, 119] and proneness to decoherence. Fig. 9(b) shows the results, displaying an increase of the average excitation number with decreasing nonharmonicity, indicating that more nonharmonic potentials are preferred when low excitation is necessitated.
@@ 4.2 Noise and decoherence
We solve noisy dynamics by averaging over different noise realizations and constructing a Monte Carlo density matrix [120]. In Figs. 9(c,d) we plot the minimal achievable infidelity at a given level of noise for the ground to the first excited state excitation protocol for a single Gaussian potential. While the positional noise affects the protocol comparably for each $\eta$, the depth noise influences the strongest the least nonharmonic case. For the specific mapping onto pointing and intensity noise in an optical tweezer setup, we have
@@
where $U_{\text{rms}}$ is a pointing root mean square (RMS) noise, and $A_{\text{rms}}$ is an intensity RMS noise. In Figs. 9(e,f) we plot the fidelity loss as a function of $U_{\text{rms}}$ and $A_{\text{rms}}$, and we find that shallower (more nonharmonic) tweezers outperform the deeper ones.

### v4

@@ 3.4 Unitaries
$\hat{P}$ is a projector onto a subspace, $\hat{U}_{\text{T}}$ is a target unitary, and $\hat{U}$ is a unitary generated through Eq. (14). Taking advantage of the optimized position of the potential $u(\tau)$ and additional slight, optimized modulation of the depth $a(\tau)$ (to increase the fidelity of the protocol), we present examples of unitaries, e.g., $\sigma_{x}$ or Hadamard operations, for the above-mentioned subspaces. In Fig. 5(a,b), we show an exemplary case of four-legged-cat basis and $\sigma_{x}$ operation performed with a Gaussian potential, where the basis states are given by
@@
$\mathcal{N}_{\pm}$ are the normalization constants, and we use $\beta=2$. Again, the protocol is performed with high fidelity, $\mathcal{F}_{\hat{U}}\approx 99\%$. In Fig. 6, we show additional examples for the Gaussian potential, including, among others, the GKP basis, where the basis is spanned by two mutually displaced GKP states,
@@
where $d_{3}$ is defined as in (27) and the squeezing and displacement need to be enough for the states not to overlap. Further unitary implementations, but for the cosine potential with the constraint (2.1), are presented in the Appendix A.
@@ 3.5 Double-well potential landscape
Figure 6: Evolution of selected orthogonal states in a Gaussian potential characterized by $\eta=0.25$ and with optimally controlled displacement $u(\tau)$ and intensity $a(\tau)$. The solid line shows $u(\tau)$ and the protocol lasts $\tau_{\text{max}}/2\pi=9$. (a-b) $\sigma_{x}$ unitary within a subspace spanned by $\psi_{0}$ and $\psi_{1}$ Fock states. The fidelity reads $\mathcal{F}_{\hat{U}}\approx 99.8\%$. (c-d) Hadamard unitary within a subspace spanned by GKP states with $r=0.7$, $(d_{1},d_{2},d_{3})=(-\sqrt{6\pi},0,\sqrt{6\pi})$. The fidelity reads $\mathcal{F}_{\hat{U}}\approx 96.4\%$. (e-f) $\sigma_{y}$ unitary within a subspace spanned by four-legged-cat states states with $\beta=2$. The fidelity reads $\mathcal{F}_{\hat{U}}\approx 96.2\%$. See the caption of Fig. 1 for subplot and curve legend details.
@@
[…]
@@
as the reward function. In Fig. 5(d,e), we show the example of a state of the two lowest vibrational levels of the left well transferred to the right one with $\mathcal{F}_{\text{st}}\approx 99\%$.
@@ 3.6 Single-shot discrimination
where $k$ is chosen such that two phase-imprinted states are nonoverlapping in phase space. After the phase imprinting, a selective measurement takes place, which can be realized in, e.g., an optical tweezer setup through the subsequent release of the trap, in which time-of-flight evolution reveals spatially separated detection clicks for each of the states [114, 44]. The reward function for such a momentum-kick protocol involves a sum of equally weighted fidelities,
[…]
Figure 8: Examples of the implementation of momentum kick protocols via an optimally controlled single-well Gaussian potential with $\eta=0.25$, $\tau_{\text{max}}/2\pi=6$, and $k=2$. (a-b) $\psi^{+}(x)=\psi_{0}(x)$ and $\psi^{-}(x)=\psi_{1}(x)$, (c-d) $\psi^{\pm}(x)=[\psi_{0}(x)\pm\psi_{1}(x)]/\sqrt{2}$, and (e-f) $\psi^{\pm}(x)=[\psi_{0}(x)\pm i\psi_{1}(x)]/\sqrt{2}$. Momentum-kick fidelities read (a-b) $\mathcal{F}_{\text{mk}}\approx 98.5\%$, (c-d) $\mathcal{F}_{\text{mk}}\approx 99.4\%$, and (e-f) $\mathcal{F}_{\text{mk}}\approx 99.5\%$. See the caption of Fig. 1 for subplot and curve legend details.
Note that in contrast to the unitary implementation, here the final relative phase between the states does not matter, as we aim only to distinguish the initial states. As an example, in Fig. 7 $\psi^{\pm}(x)$ are taken to be the two lowest eigenstates of the well, and the protocol is shown in a schematic way. In Fig. 8, we show in detail specific realizations of this and other momentum-kick protocols, utilizing a double-well Gaussian potential. They also involve other choices of orthogonal states for the discrimination, namely eigenstates of both $\sigma_{x}$ and $\sigma_{y}$ Pauli matrices. Such discrimination procedures then correspond to $\sigma_{x}$ and $\sigma_{y}$ measurements within a given two-level subspace.
The second discrimination protocol involves a double-well potential landscape and utilizes spatial separation instead of the momentum one (see Fig. 7). Initially, the states are fully confined to the left well and their discrimination is realized through a selective stealing protocol—if the state is $\psi^{+}(x)$, then it ends up in the second well after the optimal shaking, while $\psi^{-}(x)$ stays in the initial one. Then, a large spatial separation between the wells enables selective measurement. The reward function for this protocol, the selective-stealing fidelity, reads
[…]
where the definitions (42) and (43) are used. We show the exemplary case of distinguishing two lowest vibrational states in the two top panels of Fig. 7, while further examples, including discrimination of $\sigma_{x}$ and $\sigma_{y}$ eigenstates, are presented in Appendix C.
@@ 3.7 Cooling
[…]
@@
Figure 9: The accelerated algorithmic cooling method by spatial separation of the ground and excited fractions realized through optimal control of two wells. The selective-stealing cooling fidelity reads $\mathcal{F}_{\text{ss}}^{\text{cool}}\approx 97\%.$ See the caption of Fig. 1 for subplot and curve legend details.
Here, $N$ characterizes the number of excited states that are brought to the other well. The higher $N$, the more thermal state can be brought to the ground state. However, higher temperature also means a lower chance of success, as the presence of the ground state in the left well is conditioned on the subsequent destructive measurement in the right well. Hence, on average, the success rate of the protocol is proportional to the initial occupation of the ground state. Fig. 9 demonstrates this process for the four lowest states of the well, with the higher three states becoming spatially separated from the ground state. The method involving the momentum-kick protocol is analogous, but with the excited fraction being kicked out of the well. The reward function is thus
[…]
@@ 4 Feasibility
After presenting several protocols, let us now address aspects relevant to their experimental implementation. From now on, we will focus on the Fock state preparation protocol with a single-well Gaussian potential as an exemplary case. In subsection 4.1, we analyze how fast the optimized protocol can be and how excited the system is during its evolution. Then, in subsection 4.2, we check the robustness of the state preparation in the presence of noise and decoherence modeled by the stochastic fluctuations of the potential. Finally, subsection 4.3 addresses the issue of satisfying one-dimensional approximation in the specific case of an atom in an optical tweezer.
@@ 4.1 Quantum speed limit
Concerning the protocol’s speed, a fundamental restriction exists, the quantum speed limit [115], which gives the minimal time that is needed to perform a specific state transfer with a given fidelity threshold. Well-understood for two-level systems driven by static Hamiltonians [115, 116, 117], quantum speed limits have also been analyzed for time-dependent multilevel systems [118, 119], and, specifically, for the ground to the first excited state transfer in a Bose-Einstein condensate [51]. We analyze a Gaussian potential case to find a minimum time $\tau_{\text{QSL}}$ needed to perform the excitation from the ground to the first excited state, $\psi_{0}(x)\rightarrow\psi_{1}(x)$, with some infidelity threshold for different values of nonharmonicity $\eta$ . In the regime of interest, $\eta\sim 0.1-0.5$, there is no strong scaling of $\tau_{\text{QSL}}$ with the nonharmonicity and achievable fidelity scales exponentially with the protocol time. For the example of an optical tweezer, it implies that, without noise, the deeper the tweezer, the more bound states accessed and the faster the protocol (and hence requiring more control bandwidth), as $\omega\sim V_{0}^{1/2}$. We then analyze the number of excitations averaged throughout the protocol,
$\bar{n}=\int d\tau\sum_{n}n|c_{n}(\tau)|^{2},$ (53)
as this quantity is relevant for analytical quantum speed limits [117, 120, 121] and sensitivity to decoherence. Fig. 10(b) shows that the average excitation number increases as the nonharmonicity decreases. This indicates that protocols involving deeper, more harmonic potential wells tend to populate higher excited states. Interestingly, $\bar{n}$ appears to depend only weakly on $\tau_{\text{max}}$. A plausible explanation is that the maximum excitation is ultimately constrained by the finite depth of the potential and is saturated in optimally controlled scenarios. Consequently, highly nonharmonic potentials are preferred when minimizing excitation (and thus enhancing robustness) is important.
Figure 10: (a) Infidelity of the ground to the first excited state transfer, $\psi_{0}(x)\rightarrow\psi_{1}(x)$, for a Gaussian potential as a function of the length of protocol $\tau_{\text{max}}$ and nonharmonicity $\eta$. The scaling of the curve is as expected from the previous studies [51] and reveals that a quantum speed limit $\tau_{\text{QSL}}$ does not scale strongly with $\eta$. (b) Average excitation number as a function of $\tau_{\text{max}}$ and $\eta$. (c,d) The minimal achievable infidelity for given values of position ($\gamma_{1}$) and depth ($\gamma_{2}$) noises for a Gaussian potential. (e,f) The minimal achievable infidelity for given values of displacement $U_{\text{rms}}$ and intensity $A_{\text{rms}}$ noises for a rubidium-87 atom in an optical tweezer with the waist $w_{0}=710\$nm. The value of nonharmonicity $\eta$ is tuned via the tweezer’s depth $V_{0}$.
@@ 4.2 Noise and decoherence
We solve noisy dynamics by averaging over different noise realizations and constructing a Monte Carlo density matrix [122]. In Figs. 10(c,d), we plot the minimal achievable infidelity at a given level of noise for the ground to the first excited state excitation protocol for a single Gaussian potential. While the positional noise affects the protocol comparably for each $\eta$, the depth noise influences the strongest the least nonharmonic case. For the specific mapping onto pointing and intensity noise in an optical tweezer setup, we have
@@
where $U_{\text{rms}}$ is a pointing root mean square (RMS) noise, and $A_{\text{rms}}$ is an intensity RMS noise. In Figs. 10(e,f), we plot the fidelity loss as a function of $U_{\text{rms}}$ and $A_{\text{rms}}$, and we find that shallower (more nonharmonic) tweezers outperform the deeper ones.
Let us compare this observation with the charge noise in superconducting transmon qubits. In the transmon regime, the dephasing rate due to charge noise is exponentially suppressed with decreasing nonharmonicity, following $\gamma\propto\exp\small(-\sqrt{8E_{\text{J}}/E_{C}}\small)=\exp\small(-2/\eta^{2}\small)$ [28]. This trend is opposite to what we observe in Figs. 10(c,d). There, the fidelity increases with increasing $\eta$, but the dependence is at most polynomial. This suggests that, unlike optically trapped atoms, transmons benefit from less nonharmonic potentials.